OpenStudy (anonymous):

find all solutions of the system: y+x^2 = 3x y+3x = 9

OpenStudy (anonymous):

Ready to work on this together?

OpenStudy (bahrom7893):

y+3x = 9 y = 9-3x

OpenStudy (bahrom7893):

(9-3x) + x^2 =3x 9-3x+x^2-3x=0 x^2-6x+9=0

OpenStudy (bahrom7893):

(x-3)^2=0 x1=x2=3

OpenStudy (bahrom7893):

y = 9-3x = 9 - 3*3 = 0 y1=y2=0 or ur answers could just be 3 and 0

OpenStudy (bahrom7893):

or 3 and 3 and 0 and 0 depending on what class this is, i believe it's just 0 and 3

OpenStudy (bahrom7893):

Please click on Good Answer if I helped, I really want to get to the next level, thanks =)!